Blue Landscaping Inc. is preparing its budget for the first quarter of 2017. The next step in the budgeting process is to prepare a cash receipts schedule and a cash payments schedule. To that end the following information has been collected.

Clients usually pay 60% of their fee in the month that service is performed, 30% the month after, and 10% the second month after receiving service.

Actual service revenue for 2016 and expected service revenues for 2017 are November 2016, $92,930; December 2016, $81,950; January 2017, $102,750; February 2017, $124,450; March 2017, $134,080.

Purchases of landscaping supplies (direct materials) are paid 60% in the month of purchase and 40% the following month. Actual purchases for 2016 and expected purchases for 2017 are December 2016, $16,880; January 2017, $16,090; February 2017, $19,480; March 2017, $23,120.Prepare the following schedules for each month in the first quarter of 2017 and for the quarter in total: (1) Expected collec
Answer #1

Answer-1)-

BLUE LANDESCAPING INC.
SCHEDULE OF EXPECTED COLLECTIONS FROM CIENTS
FOR THE FIRST QUARTER OF 2017
PARTICULARS JANUARY FEBRUARY MARCH QUARTER
$ $ $ $
November $92930*10%= $9293 9293
December $81950*30%= $24585 $81950*10%= $8195 32780
January $102750*60%= $61650 $102750*30%= $30825 $102750*10%= $10275 102750
Febuary $124450*60%= $74670 $124450*30%= $37335 112005
March $134080*60%= $80448 80448
Total collections $ 95528 113690 128058 337276

2)-

BLUE LANDESCAPING INC.
SCHEDULE OF EXPECTED PAYMENTS FOR LANDSCAPING SUPPLIES
FOR THE FIRST QUARTER OF 2017
PARTICULARS JANUARY FEBRUARY MARCH QUARTER
$ $ $ $
December $16880*40%= $6752 6752
January $16090*60%= $9654 $16090*40%= $6436 16090
Febuary $19480*60%= $11688 $19480*40%= $7792 19480
March $23120*60%= $13872 13872
Total payments $ 16406 18124 21664 56194